7x^2-y-2x=0 solve for y.
Found 2 solutions by checkley75, pleasuretohelp:
Answer by checkley75(3666)

7X^2-Y-2X=0
-Y=-7X^2+2X
Y=7X^2-2X
Y=X(7X-2)

Answer by pleasuretohelp(9)

7x^2-y-2x=0
First, move all components that do not have y to the right side.
7x^2 - y - 2x = 0
- y = -7x^2 + 2x
Second, divide both sides by -1.
- y/-1 = -7x^2/-1 + 2x/-1
y = 7x^2 - 2x